OpenStudy (anonymous):

Three friends entered a contest to count the number of jelly beans in a jar. The table shows their guesses of the number of jelly beans in the jar. There were actually 950 jelly beans in the jar. Which statements are true? Name Number guessed Brandon 991 Li Mei 939 Jodi 976 Choose exactly two answers that are correct. A. Li Mei’s estimate had about a 1.2% error. B. Jodi had the highest percent error. C. Jodi’s estimate had about a 26% error. D. Brandon’s estimate had about a 4.3% error.

OpenStudy (anonymous):

@jamaica25

OpenStudy (texaschic101):

total of 950 jb's Brandon -- (991 - 950) / 950 = 41/950 = 0.043 = 4.3% error Li Mei -- (950 - 939) / 939 = 11/950 = 0.0115 = 1.1% error Jody -- (976 - 950) / 950 = 26/950 = 0.027 = 2.7% error I would have to go with A and D